# PromptAD: Learning Prompts with only Normal Samples for Few-Shot Anomaly Detection

### v1

@@ 4.2 Semantic Concatenation
where z denotes normal visual feature, $\bar{\textbf{w}}^{n}=\frac{\sum_{i=1}^{N}g(\textbf{s}^{n}_{i})}{N}$ is the prototype of normal prompt features, $\mathcal{W}=\{g(\textbf{s})|\textbf{s}\in\textup{MAPs}\cup\textup{LAPs}\}$ is a set containing all anomaly prompt features. Since more negative samples can produce a better contrastive learning effect [18], each anomaly prompt feature is compared with the visual feature
@@ G Visualization Results of Attention Map
Bottle 98.7$\pm$0.6 97.4$\pm$0.7 99.4$\pm$0.4 98.2$\pm$0.9 99.8$\pm$0.3 99.5$\pm$0.1 98.5$\pm$1.0 99.2$\pm$0.3 99.3$\pm$0.3 100.0$\pm$0.1 99.5$\pm$0.2 98.8$\pm$0.2 99.2$\pm$0.3 99.3$\pm$0.4 100.0$\pm$0.2
Cable 71.2$\pm$3.3 57.7$\pm$4.6 88.8$\pm$4.2 88.9$\pm$1.9 94.2$\pm$1.2 76.2$\pm$5.2 62.3$\pm$5.9 91.0$\pm$2.7 88.4$\pm$0.7 99.9$\pm$1.5 83.4$\pm$3.1 70.0$\pm$6.1 91.0$\pm$2.7 90.9$\pm$0.9 98.8$\pm$0.9
Capsule 70.2$\pm$3.0 57.7$\pm$7.3 67.8$\pm$2.9 72.3$\pm$6.8 84.6$\pm$6.7 70.9$\pm$6.1 64.3$\pm$3.0 72.8$\pm$7.0 77.3$\pm$8.8 100.0$\pm$7.7 78.9$\pm$5.5 65.2$\pm$2.5 72.8$\pm$7.0 82.3$\pm$8.9 100.0$\pm$1.5
Garpet 98.1$\pm$0.2 96.6$\pm$1.0 95.3$\pm$0.8 99.8$\pm$0.3 100.0$\pm$0.0 98.3$\pm$0.4 97.8$\pm$0.5 96.6$\pm$0.5 99.8$\pm$0.3 100.0$\pm$0.0 98.6$\pm$0.2 97.9$\pm$0.4 96.6$\pm$0.5 100.0$\pm$0.0 100.0$\pm$0.0
Grid 40.0$\pm$6.8 54.2$\pm$6.7 63.6$\pm$10.3 99.5$\pm$0.3 99.8$\pm$0.9 41.3$\pm$3.6 67.2$\pm$4.2 67.7$\pm$8.3 99.4$\pm$0.2 98.7$\pm$1.2 44.6$\pm$6.6 68.1$\pm$3.8 67.7$\pm$8.3 99.6$\pm$0.1 98.6$\pm$0.5
Hazelnut 95.8$\pm$1.3 88.3$\pm$2.6 88.3$\pm$2.7 97.5$\pm$1.4 99.8$\pm$0.8 96.2$\pm$2.1 90.8$\pm$0.8 93.2$\pm$3.8 98.3$\pm$0.7 99.4$\pm$0.6 98.4$\pm$1.3 91.9$\pm$1.2 93.2$\pm$3.8 98.4$\pm$0.4 99.8$\pm$0.2
Leather 100.0$\pm$0.0 97.5$\pm$0.7 97.3$\pm$0.7 99.9$\pm$0.0 100.0$\pm$0.0 100.0$\pm$0.0 97.5$\pm$0.9 97.9$\pm$0.7 99.9$\pm$0.0 91.8$\pm$0.1 100.0$\pm$0.0 98.5$\pm$0.2 97.9$\pm$0.7 100.0$\pm$0.0 95.4$\pm$0.1
Metal nut 71.0$\pm$2.2 53.0$\pm$3.8 73.4$\pm$2.9 98.7$\pm$0.8 99.1$\pm$0.5 77.0$\pm$7.9 54.8$\pm$3.8 77.7$\pm$8.5 99.4$\pm$0.2 91.3$\pm$0.7 77.8$\pm$5.7 60.7$\pm$5.2 77.7$\pm$8.5 99.5$\pm$0.2 91.5$\pm$0.1
Pill 86.5$\pm$3.1 61.3$\pm$3.8 81.9$\pm$2.8 91.2$\pm$2.1 92.6$\pm$1.5 84.8$\pm$0.9 59.1$\pm$6.4 82.9$\pm$2.9 92.3$\pm$0.7 100.0$\pm$0.8 86.7$\pm$0.3 54.9$\pm$2.7 82.9$\pm$2.9 92.8$\pm$1.0 99.8$\pm$0.8
Screw 46.7$\pm$2.5 55.0$\pm$2.5 44.4$\pm$4.6 86.4$\pm$0.9 65.0$\pm$2.9 46.6$\pm$2.2 54.0$\pm$4.4 49.0$\pm$3.8 86.0$\pm$2.1 98.6$\pm$2.0 50.5$\pm$5.4 50.0$\pm$4.1 49.0$\pm$3.8 87.9$\pm$1.2 100.0$\pm$2.3
Tile 99.9$\pm$0.1 92.2$\pm$2.2 99.0$\pm$0.9 99.9$\pm$0.0 100.0$\pm$0.0 99.9$\pm$0.1 93.3$\pm$1.1 98.5$\pm$1.0 99.9$\pm$0.2 93.6$\pm$0.1 100.0$\pm$0.0 93.1$\pm$0.6 98.5$\pm$1.0 99.9$\pm$0.1 92.9$\pm$0.1
Toothbursh 71.7$\pm$2.6 82.5$\pm$1.2 83.3$\pm$3.8 92.2$\pm$4.9 98.9$\pm$1.0 78.6$\pm$3.2 87.6$\pm$4.2 85.9$\pm$3.5 97.5$\pm$1.6 71.0$\pm$1.5 78.8$\pm$5.2 89.2$\pm$2.5 85.9$\pm$3.5 96.7$\pm$2.6 83.6$\pm$0.6
Transistor 77.2$\pm$2.0 73.3$\pm$6.0 78.1$\pm$6.9 83.4$\pm$3.8 94.0$\pm$6.5 81.3$\pm$3.7 72.8$\pm$6.3 90.0$\pm$4.3 85.3$\pm$1.7 97.5$\pm$4.0 81.4$\pm$2.1 82.4$\pm$6.5 90.0$\pm$4.3 85.7$\pm$2.5 98.1$\pm$3.8
Wood 98.8$\pm$0.3 96.1$\pm$1.2 97.8$\pm$0.3 99.9$\pm$0.1 97.9$\pm$0.3 99.2$\pm$0.4 96.9$\pm$0.5 98.3$\pm$0.6 99.9$\pm$0.1 97.4$\pm$0.1 98.9$\pm$0.6 97.0$\pm$0.2 98.3$\pm$0.6 99.8$\pm$0.3 95.6$\pm$0.3

### v2

@@ 4.2 Semantic Concatenation
where z denotes normal visual feature, $\bar{\textbf{w}}^{n}=\frac{\sum_{i=1}^{N}g(\textbf{s}^{n}_{i})}{N}$ is the prototype of normal prompt features, $\mathcal{W}=\{g(\textbf{s})|\textbf{s}\in\textup{MAPs}\cup\textup{LAPs}\}$ is a set containing all anomaly prompt features. Since more negative samples can produce a better contrastive learning effect [18], each anomaly prompt feature is compared with the visual feature.
@@ G Visualization Results of Attention Map
Bottle 98.7$\pm$0.6 97.4$\pm$0.7 99.4$\pm$0.4 98.2$\pm$0.9 99.8$\pm$0.3 99.5$\pm$0.1 98.5$\pm$1.0 99.2$\pm$0.3 99.3$\pm$0.3 99.4$\pm$0.1 99.5$\pm$0.2 98.8$\pm$0.2 99.2$\pm$0.3 99.3$\pm$0.4 99.8$\pm$0.2
Cable 71.2$\pm$3.3 57.7$\pm$4.6 88.8$\pm$4.2 88.9$\pm$1.9 94.2$\pm$1.2 76.2$\pm$5.2 62.3$\pm$5.9 91.0$\pm$2.7 88.4$\pm$0.7 91.8$\pm$1.5 83.4$\pm$3.1 70.0$\pm$6.1 91.0$\pm$2.7 90.9$\pm$0.9 95.4$\pm$0.9
Capsule 70.2$\pm$3.0 57.7$\pm$7.3 67.8$\pm$2.9 72.3$\pm$6.8 84.6$\pm$6.7 70.9$\pm$6.1 64.3$\pm$3.0 72.8$\pm$7.0 77.3$\pm$8.8 91.3$\pm$7.7 78.9$\pm$5.5 65.2$\pm$2.5 72.8$\pm$7.0 82.3$\pm$8.9 91.5$\pm$1.5
Carpet 98.1$\pm$0.2 96.6$\pm$1.0 95.3$\pm$0.8 99.8$\pm$0.3 100.0$\pm$0.0 98.3$\pm$0.4 97.8$\pm$0.5 96.6$\pm$0.5 99.8$\pm$0.3 100.0$\pm$0.0 98.6$\pm$0.2 97.9$\pm$0.4 96.6$\pm$0.5 100.0$\pm$0.0 100.0$\pm$0.0
Grid 40.0$\pm$6.8 54.2$\pm$6.7 63.6$\pm$10.3 99.5$\pm$0.3 99.8$\pm$0.9 41.3$\pm$3.6 67.2$\pm$4.2 67.7$\pm$8.3 99.4$\pm$0.2 99.9$\pm$1.2 44.6$\pm$6.6 68.1$\pm$3.8 67.7$\pm$8.3 99.6$\pm$0.1 98.8$\pm$0.5
Hazelnut 95.8$\pm$1.3 88.3$\pm$2.6 88.3$\pm$2.7 97.5$\pm$1.4 99.8$\pm$0.8 96.2$\pm$2.1 90.8$\pm$0.8 93.2$\pm$3.8 98.3$\pm$0.7 100.0$\pm$0.6 98.4$\pm$1.3 91.9$\pm$1.2 93.2$\pm$3.8 98.4$\pm$0.4 99.8$\pm$0.2
Leather 100.0$\pm$0.0 97.5$\pm$0.7 97.3$\pm$0.7 99.9$\pm$0.0 100.0$\pm$0.0 100.0$\pm$0.0 97.5$\pm$0.9 97.9$\pm$0.7 99.9$\pm$0.0 100.0$\pm$0.1 100.0$\pm$0.0 98.5$\pm$0.2 97.9$\pm$0.7 100.0$\pm$0.0 100.0$\pm$0.1
Metal nut 71.0$\pm$2.2 53.0$\pm$3.8 73.4$\pm$2.9 98.7$\pm$0.8 99.1$\pm$0.5 77.0$\pm$7.9 54.8$\pm$3.8 77.7$\pm$8.5 99.4$\pm$0.2 98.6$\pm$0.7 77.8$\pm$5.7 60.7$\pm$5.2 77.7$\pm$8.5 99.5$\pm$0.2 100.0$\pm$0.1
Pill 86.5$\pm$3.1 61.3$\pm$3.8 81.9$\pm$2.8 91.2$\pm$2.1 92.6$\pm$1.5 84.8$\pm$0.9 59.1$\pm$6.4 82.9$\pm$2.9 92.3$\pm$0.7 93.6$\pm$0.8 86.7$\pm$0.3 54.9$\pm$2.7 82.9$\pm$2.9 92.8$\pm$1.0 92.9$\pm$0.8
Screw 46.7$\pm$2.5 55.0$\pm$2.5 44.4$\pm$4.6 86.4$\pm$0.9 65.0$\pm$2.9 46.6$\pm$2.2 54.0$\pm$4.4 49.0$\pm$3.8 86.0$\pm$2.1 71.0$\pm$2.0 50.5$\pm$5.4 50.0$\pm$4.1 49.0$\pm$3.8 87.9$\pm$1.2 83.6$\pm$2.3
Tile 99.9$\pm$0.1 92.2$\pm$2.2 99.0$\pm$0.9 99.9$\pm$0.0 100.0$\pm$0.0 99.9$\pm$0.1 93.3$\pm$1.1 98.5$\pm$1.0 99.9$\pm$0.2 100.0$\pm$0.1 100.0$\pm$0.0 93.1$\pm$0.6 98.5$\pm$1.0 99.9$\pm$0.1 100.0$\pm$0.1
Toothbrush 71.7$\pm$2.6 82.5$\pm$1.2 83.3$\pm$3.8 92.2$\pm$4.9 98.9$\pm$1.0 78.6$\pm$3.2 87.6$\pm$4.2 85.9$\pm$3.5 97.5$\pm$1.6 97.5$\pm$1.5 78.8$\pm$5.2 89.2$\pm$2.5 85.9$\pm$3.5 96.7$\pm$2.6 98.1$\pm$0.6
Transistor 77.2$\pm$2.0 73.3$\pm$6.0 78.1$\pm$6.9 83.4$\pm$3.8 94.0$\pm$6.5 81.3$\pm$3.7 72.8$\pm$6.3 90.0$\pm$4.3 85.3$\pm$1.7 97.4$\pm$4.0 81.4$\pm$2.1 82.4$\pm$6.5 90.0$\pm$4.3 85.7$\pm$2.5 95.6$\pm$3.8
Wood 98.8$\pm$0.3 96.1$\pm$1.2 97.8$\pm$0.3 99.9$\pm$0.1 97.9$\pm$0.3 99.2$\pm$0.4 96.9$\pm$0.5 98.3$\pm$0.6 99.9$\pm$0.1 98.7$\pm$0.1 98.9$\pm$0.6 97.0$\pm$0.2 98.3$\pm$0.6 99.8$\pm$0.3 98.6$\pm$0.3